Solve the problem.
Police use the formula s = square root of 30fd to estimate the speed of s of a a car in miles per hour, where d is the distance in feet that the car skidded and f is the coefficient of friction. If the coefficient of friction on a certain gravel road is 0.23 and a car skidded 330 feet, find the speed of the car, to the nearest mile per hour.

Answer by richwmiller(17219)

s=sqrt(30fd)
s=sqrt(30*.23*330)
s=sqrt(2277)
s=47.717=48 mph